Find A and write the following linear system in the matrix equation form (λI2 − A)X = 0


−x + y = λx

3x + y = λy


For the systems above, find:


(10.1) The determinant (known as the characteristic equation).

(10.2) Solve for λ when det(λI2 − A) = 0.

(10.3) Substitute for each value of λ from (ii) into the equation (λI2 − A)X = 0 and solve the corresponding system for X = x

y


1
Expert's answer
2021-05-28T08:18:15-0400

Rewriting the equations,

"x(-1-\lambda)+y=0"

"3x+y(1-\lambda)=0"


(1) "\begin{bmatrix}
 (-1-\lambda) & 1 \\
 3 & (1-\lambda)
\end{bmatrix}\begin{bmatrix}
 x \\
 y
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0
\end{bmatrix}"


"(-1-\lambda)(1-\lambda)-3=-1+\lambda-\lambda+\lambda^2-3"

"=\lambda^2-4"


(2) when "\det(\lambda I_2-A)=0"

"\Rightarrow\lambda^2-4=0"

"\Rightarrow\lambda=\pm2"


(3) Substituting the value of "(\lambda I_2-A)X=0"

For system to be consistent, "\lambda\cancel=\pm2"

"(\lambda I_2-A)X=0"

"\begin{bmatrix}
 (-1-\lambda) & 1 \\
 3 & (1-\lambda)
\end{bmatrix}\begin{bmatrix}
 x \\
 y
\end{bmatrix}=\begin{bmatrix}
 0 \\
 0
\end{bmatrix}"


"x(-1-\lambda)+y=0"

"x(-1-2)+y=0"

"-3x+y=0"

"3x-y=0"


"3x+y(1-\lambda)=0"

"3x+y(1-2)=0"

"3x-y=0"


Since both the equations are parallel therefore the system has infinitely many solutions.